a typical jar that has been tightened to a reasonable degree requires 2.0 N•m to open. if you grab a 7.0cm diameter jar lid with one hand so that your thumb and fingers exert equal magnitude forces on opposite sides of the lid, what is the magnitude of each of the forces?
Answer #1

required torque to open the lid = 2 N.m

since finger pushes like this  

so

0.07*F = 2

F = 2/0.07 =28.5714286 N

so force of 28.5714286 N needs to be supplied by finger and thumb to open this jar

Good Luck !! for exam
let me know in a comment if there is any problem or doubts please wait for me to respond.